Candice has a 76 average in Physics class going into the final, and that final counts for 40% of the course grade. What must she score on the final to get an 80 in the class?

Answer by Edwin McCravy(20056)

There are several ways to look at it.  Here's one that seems to fit in
with the way people think nowadays:

On a scale of 1 to 10, 

1. the average up to the final, 76, gets a 6, so multiply it by 6.

2. the final exam, X, gets a 4, so multiply it by 4.

3. the course grade desired, 80, gets a 10, so multiply it by 10.

(76)(6) + (4)(X) = (80)(10)

456 + 4X = 800
      4X = 344
       X = 86

Edwin